# Interacting tachyon with varying mass dark matter

### v1

@@ III Phase space analysis of autonomous system (24):
The equation (24) reduces to
$\frac{dx}{dN}$ $=-3x-\beta z+6xy+3xz+2(\alpha+\beta)zy+\text{higher degree terms}.$ (58)
$\frac{dy}{dN}$ $=-3y-\frac{3}{2}y^{2}-\frac{\beta}{2}xz+3zy+\text{higher degree terms}.$ (59)
$\frac{dz}{dN}$ $=3zy-\frac{15}{2}zy^{2}-3z^{2}y+\frac{3}{4}zx^{2}+\text{higher degree terms}.$ (60)
Now for the variables $(X,Y,Z)$ the system (58-60) takes the following form:
$\frac{dX}{dN}$ $=-3X+6X^{2}-\frac{9}{\beta}Z^{2}+6XY+3\left(1+\frac{3+2\alpha}{\beta}\right)ZX+3\left(\frac{\beta+2\alpha}{\beta}\right)ZY+\text{higher degree terms}.$ (61)
$\frac{dY}{dN}$ $=-3Y-\left(\frac{15}{2}+\beta\right)X^{2}+\frac{3}{2}Y^{2}+\left(\frac{3}{2}-\frac{9}{\beta}\right)Z^{2}-\left(9+\beta\right)XY+\left(\beta-\frac{6\alpha}{\beta}+\frac{3}{2}\right)ZX$ (62)
$+\left(\beta+\frac{9-6\alpha}{\beta}\right)ZY+\text{higher degreeterms}.$ (63)
$\frac{dZ}{dN}$ $=3Z\left(X+Y\right)+\text{higher degree terms}.$ (64)
The center manifold at $C$ is obtained by
$X$ $=$ $-\frac{3}{\beta}Z^{2}+\text{higher degree terms}.$ (65)
$Y$ $=$ $(\frac{3}{\beta}+\frac{1}{2})Z^{2}+\text{higher degree terms}.$ (66)
Then the flow along the center manifold can be written as
$\frac{dZ}{dN}=\frac{3}{2}Z^{3}+\text{higher degree terms}.$ (68)
@@
The vector fields of (57) and (68) are topologically conjugate near the critical points $B$ and $C$ respectively of the system (24). So the flows are unstable on the center manifold near the $B$ and $C$. Thus the critical points $B$ and $C$ are saddle-node in nature. Hence, if the initial state is on the $xy$ plane we have late time de-Sitter solution.
@@
$\bullet$ The critical points $F$ and $G$ exists for all $\alpha,~{}\beta>0$. If we fluctuate $\alpha$ or $\beta$ we get non-isolated critical points $F$ and $G$. These points correspond to the scaling solutions and the cosmic coincidence problem can be alleviated depending on the values of $\alpha$ and $\beta$. Near these critical points dark matter and dark energy correspond to the form of dust and the evolution of the universe behaves as quintessence boundary for $2\alpha=\beta$. On the other hand, there exists an accelerating universe when $2\alpha>\beta$. The critical points $F$ and $G$ are situated on the $z=1$ plane and non-hyperbolic in nature. As the real parts of all eigenvalues are zero, it is very difficult to find the character of the vector field around these points analytically. We depict phase portrait numerically (see FIGs. 4 and 4) and try to understand the nature of flow near these points. Now it is to be noted from our numerical analysis that $z=1$ plane is an attractor. The vector field is divided into two chambers by $xz$ plane. If the initial position of the flow is in the positive (negative) $y$ coordinate but not on $z=1$ plane then the flow approaches towards $F$ (G). But as soon as the flow touches the $z=1$ plane, it behaves as outgoing spiral. The vector field flows anticlockwise along out going spiral around the critical point $F$ and clockwise around $G$ when we see from the top of $z=1$ plane (see FIGs. 4 and 4).
@@
Figure 5: The figures show the evolution of cosmological parameters numerically for different choices of initial conditions and $\alpha=\beta=1$. In panel (a) and (b), the starting position of the trajectory is near the critical point $B$ or $C$. In this case, we get scalar field dominated late-time solution. The cosmic evolution near the points characterizes the de-Sitter expansion of the universe. In panel (c) and (d), the starting position of the trajectory is near the critical point $D$ or $E$. There is a brief period of matter dominated deceleration expansion phase of the Universe before the transition to the scalar field dominated de-Sitter expansion of the universe. In panel (e) and (f), the starting position of the trajectory is near the critical point $F$ or $G$. As the trajectory goes towards $z=1$ plane in a spiral-like manner (see fig (4)), the values of cosmological parameters oscillate between a particular range. The nearer the trajectory goes towards $z=1$ plane the amplitude of oscillation decreases and the values of deceleration parameter oscillate between $-0.18$ to $-0.38$ whereas the values of effective equation of state parameter oscillate between $-0.4$ to $-0.6$. On the other hand, the density parameters oscillate between $0.4$ to $0.6$ (a scaling solution) near the $z=1$ plane.
@@ IV Statefinder Diagnosis
in which the over dot and $q=-\ddot{a}/aH^{2}$ denote the differentiation with respect to the cosmic time $t$ and deceleration parameter, respectively. Different combinations of the pair ($r,s$) could serve as distinctive representation of different DE models. If ($r=1,s=0$), then the DE behaves like cosmological constant (i.e., $\Lambda CDM$ model). On the other hand, ($r<1,s>0$) (or, $r=1,s=1$) suggests that the DE is a Quintessence (or, SCDM). For the case of Chaplygin gas model, the trajectories in the $s-r$ plane lie in the region where $s<0$ and $r>1$ [46]. This makes it a powerful tool for effectively distinguishing between various DE models, even when they produce similar expansion histories. However, one can look into the references [47, 48, 49, 50, 51, 52, 53, 54, 55, 56], where a detailed statefinder pair analysis for different DE models is comprehensively discussed.
In this work, the present model is scrutinized employing statefinder diagnostic tools. The process involves calculating statefinder parameters with respect to redshift, compare them with that of $\Lambda CDM$, and exploring their behavior at both high and low redshift limits. Using the energy conservation equations (15) and (16), we obtain (after some algebraic calculations and assumptions) the energy densities of DM and DE as
@@
$\centering E(z)=\frac{H}{H_{0}}=(1+z)^{\frac{3}{2}\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}+1\right)}\left[\Omega_{m0}+\Omega_{\phi 0}~{}(1+z)^{-3\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}-\omega_{\phi}\right)}\left\{1+\frac{\Omega_{m0}\alpha\sqrt{3(1+\omega_{\phi})}}{3\sqrt{3}\Omega_{\phi 0}\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}-\omega_{\phi}\right)}(1+z)^{-3\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}+1\right)}\right\}\right]^{\frac{1}{2}},\@add@centering$ (76)
@@
$r(z)=\frac{1}{2}\left(\frac{\Omega_{m0}(z+1)^{-2\alpha\sqrt{\omega_{\phi}+1}+3\omega_{\phi}-3}\left(\Omega_{\phi 0}(z+1)^{\alpha\left(-\sqrt{\omega_{\phi}+1}\right)}\left(9\omega_{\phi}^{2}\Psi\left(3\omega_{\phi}-\alpha\left(3\sqrt{\omega_{\phi}+1}-\alpha\right)\right)+\Theta\right)+\chi\right)}{\Xi^{2}\left(3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}\right)}+\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+1\right)\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+2\right)\right),$ (78)
@@
$s(z)=\frac{\frac{\Omega_{m0}(z+1)^{-2\alpha\sqrt{\omega_{\phi}+1}+3\omega_{\phi}-3}\left(\Omega_{\phi 0}(z+1)^{\alpha\left(-\sqrt{\omega_{\phi}+1}\right)}\left(9\omega_{\phi}^{2}\Psi\left(3\omega_{\phi}-\alpha\left(3\sqrt{\omega_{\phi}+1}-\alpha\right)\right)+\Theta\right)+\chi\right)}{\Xi^{2}\left(3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}\right)}+\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+1\right)\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+2\right)-2}{3\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}\right)}$ (79)
@@
$\Delta=(z+1)^{-2\alpha\sqrt{\omega_{\phi}+1}+3\omega_{\phi}-3}\left(-\Omega_{m0}\alpha\sqrt{\omega_{\phi}+1}+\frac{\Omega_{m0}\alpha\left(\alpha+\alpha\omega_{\phi}+3\sqrt{\omega_{\phi}+1}\right)}{3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}}+\Omega_{\phi 0}\left(3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}\right)(z+1)^{\alpha\sqrt{\omega_{\phi}+1}+3}\right)$,
@@
$\chi=-\Omega_{m0}\alpha\left(3\omega_{\phi}^{2}\left(3\sqrt{\omega_{\phi}+1}-4\alpha\right)+2\left(2\alpha^{2}-9\right)\sqrt{\omega_{\phi}+1}\omega_{\phi}+4\alpha\left(\alpha\sqrt{\omega_{\phi}+1}+3\right)+9\sqrt{\omega_{\phi}+1}\right)$
and
@@
${}~{}~{}~{}~{}~{}-\alpha\left(\alpha\left(\alpha\sqrt{\omega_{\phi}+1}((1+z)^{3\omega_{\phi}}+\Psi)+6(1+z)^{3\omega_{\phi}}\right)+9\sqrt{\omega_{\phi}+1}(1+z)^{3\omega_{\phi}}\right)$.
The evolutionary trajectories of the statefinder pair in the $s$-$r$ plane are shown in figures 6 for different choices of the model parameter $\alpha$ and $\omega_{\phi}=-0.95$. The evolutionary trajectories of statefinder pair $(s,r)$ of our model start its evolution along the direction of $s$ increasing and $r$ decreasing and with time, trajectories pass through the $\Lambda CDM$ fixed point $(s=0,r=1)$. After making a swirl, these lies in the region $(s<0,r>1)$ in the future. Also for different values of the model parameter $\alpha$, we found different trajectories and present values of the statefinder pair $(s_{0},r_{0})$ as shown by the colored dots. This scenario shows that $\alpha$ affects on the evolutionary trajectories in the $s$-$r$ plane. Moreover, in figure 7, we obtain same evolutionary curve in $s$-$r$ plane for different values of the model parameter $\alpha$ and $\omega_{\phi}=-1$. In contrary to figure 6, we found that $\alpha$ does not affect on the evolutionary trajectories in the $s$-$r$ plane for the case $\omega_{\phi}=-1$. Furthermore, we have also shown the evolutionary trajectories of another statefinder pair $(q,r)$ for the model in figures 6 and 7 . From figure 6, we found that for $\alpha=0.000001$, the evolutionary curve of statefinder pair $(q,r)$ of the model starts from the $SCDM$ $(r=1,q=0.5)$ in the past while the other evolution curves for the model have a smaller deviation from this fixed point and finally all the curves reach above the de Sitter expansion (SS) $(r=1,q=-1)$ in the future. We also found from figure 7 that $\alpha$ has no effect on the evolutionary trajectories in the $q$-$r$ plane for $\omega_{\phi}=-1$. Moreover, figures 6, 7 clearly indicate that the evolution of $q$ shows a smooth signature flip from its positive value regime to negative value regime in the $q$-$r$ plane. Hence, from the statefinder diagnostics analysis, the present model explains the late-time accelerated universe and also the transition from the early decelerated phase ($q>0$) to the current accelerated phase ($q<0$).

### v2

@@ III Phase space analysis of autonomous system (LABEL:autonomous_system_2):
The equation (LABEL:autonomous_system_2) reduces to
$\frac{dx}{dN}$ $=\left(1-x^{2}\right)\left(\alpha\sqrt{1-x^{2}}z+3x(y-1)^{2}(z-1)-(y-1)^{2}z(\alpha+\beta)\right).$ (45)
$\frac{dy}{dN}$ $=\frac{1}{2}(y-1)^{3}\left(3(1-z)\left(1-\sqrt{1-x^{2}}(y-1)^{2}\right)+\beta xz\right).$ (46)
$\frac{dz}{dN}$ $=\frac{3}{2}(y-1)^{2}(1-z)^{2}z\left(1-\sqrt{1-x^{2}}(y-1)^{2}\right).$ (47)
At the origin the jacobian matrix takes the following form
$J(0,0,0)=\begin{bmatrix}-3&0&-\beta\\ 0&-3&0\\ 0&0&0\end{bmatrix}.$
Similar to the case of critical point $B$, we need the same $T$ to diagonalize $J(0,0,0)$ which is given by
$T=\begin{bmatrix}1&0&-\frac{\beta}{3}\\ 0&1&0\\ 0&0&1\end{bmatrix}$
By the orientation preserving homeomorphism $T$ (as $\det(T)>0$), we get the new $(P,Q,R)$-coordinate system and the autonomous equations (45-47) take the following form:
$\frac{dP}{dN}=\frac{1}{2}\beta(Q-1)^{2}(R-1)^{2}R\left(1-(Q-1)^{2}\sqrt{1-\left(P-\frac{\beta R}{3}\right)^{2}}\right)\\ \frac{1}{27}\left((\beta R-3P)^{2}-9\right)\left(9P(Q-1)^{2}(R-1)+\alpha R\sqrt{9-(\beta R-3P)^{2}}-3(Q-1)^{2}R(\alpha+\beta R)\right).$ (48)
$\frac{dQ}{dN}$ $=\frac{1}{2}(Q-1)^{3}\left(3(1-R)\left(1-(Q-1)^{2}\sqrt{1-\left(P-\frac{\beta R}{3}\right)^{2}}\right)+\beta R\left(P-\frac{\beta R}{3}\right)\right).$ (49)
$\frac{dR}{dN}$ $=\frac{3}{2}(Q-1)^{2}(R-1)^{2}R\left(1-(Q-1)^{2}\sqrt{1-\left(P-\frac{\beta R}{3}\right)^{2}}\right).$ (50)
And ignoring the higher degree terms in the equations (48-50) one can obtain following simplified form
$\frac{dP}{dN}$ $=-3P+3PR+6PQ+(2\alpha+\beta)QR+\text{~{}higher degree terms}.$ (51)
$\frac{dQ}{dN}$ $=-3Q+\frac{21}{2}Q^{2}+3QR-\frac{\beta}{2}PR+\text{~{}higher degree terms}.$ (52)
$\frac{dR}{dN}$ $=-6R^{2}Q-\frac{15}{2}Q^{2}R+\text{~{}higher degree terms}.$ (53)
Then using center manifold theory we able to conclude that the flow along the center manifold at $C$ is topologically equivalent to the flow of
$\frac{dR}{dN}\approx 0.$ (54)
@@
The vector fields of (44) and (54) are topologically conjugate near the critical points $B$ and $C$ respectively of the system (LABEL:autonomous_system_2). The flow is almost static along the center manifold. As a result of this, the vector field is dominated by the flow on $PR$-plane. We also note that the nature of vector field does not depend on $\beta$ and the vector fields are stable near the critical points $B$ and $C$. Thus the critical points $B$ and $C$ are stable-node in nature and if the initial state is near $B$ or $C$ we have late time de-Sitter solution.
@@
$\bullet$ The critical points $F$ and $G$ exists for all $\alpha,~{}\beta>0$. If we fluctuate $\alpha$ or $\beta$ we get non-isolated critical points $F$ and $G$. These points correspond to the scaling solutions and the cosmic coincidence problem can be alleviated depending on the values of $\alpha$ and $\beta$. Near these critical points dark matter and dark energy correspond to the form of dust and the evolution of the universe behaves as quintessence boundary for $2\alpha=\beta$. On the other hand, there exists an accelerating universe when $2\alpha>\beta$. The critical points $F$ and $G$ are situated on the $z=1$ plane and non-hyperbolic in nature. As the real parts of all eigenvalues are zero, it is very difficult to find the character of the vector field around these points analytically. We depict phase portrait numerically (see figures 5 and 5) and try to understand the nature of flow near these points. Now it is to be noted from our numerical analysis that $z=1$ plane is an attractor. The vector field is divided into two chambers by $xz$ plane. If the initial position of the flow is in the positive (negative) $y$ coordinate but not on $z=1$ plane then the flow approaches towards $F$ (G). But as soon as the flow touches the $z=1$ plane, it behaves as outgoing spiral. The vector field flows anticlockwise along out going spiral around the critical point $F$ and clockwise around $G$ when we see from the top of $z=1$ plane (see figures 5 and 5).
@@
Figure 6: The figures show the evolution of cosmological parameters numerically for different choices of initial conditions and $\alpha=\beta=1$. In panel (a) and (b), the starting position of the trajectory is near the critical point $B$ or $C$. In this case, we get scalar field dominated late-time solution. The cosmic evolution near the points characterizes the de-Sitter expansion of the universe. In panel (c) and (d), the starting position of the trajectory is near the critical point $D$ or $E$. There is a brief period of matter dominated deceleration expansion phase of the Universe before the transition to the scalar field dominated de-Sitter expansion of the universe. In panel (e) and (f), the starting position of the trajectory is near the critical point $F$ or $G$. As the trajectory goes towards $z=1$ plane in a spiral-like manner (see figure 5), the values of cosmological parameters oscillate between a particular range. The nearer the trajectory goes towards $z=1$ plane the amplitude of oscillation decreases and the values of deceleration parameter oscillate between $-0.18$ to $-0.38$ whereas the values of effective equation of state parameter oscillate between $-0.4$ to $-0.6$. On the other hand, the density parameters oscillate between $0.4$ to $0.6$ (a scaling solution) near the $z=1$ plane.
@@ IV Statefinder Diagnosis
in which the over dot and $q=-\ddot{a}/aH^{2}$ denote the differentiation with respect to the cosmic time $t$ and deceleration parameter, respectively. Different combinations of the pair ($r,s$) could serve as distinctive representation of different DE models. If ($r=1,s=0$), then the DE behaves like cosmological constant (i.e., $\Lambda CDM$ model). On the other hand, ($r<1,s>0$) (or, $r=1,s=1$) suggests that the DE is a Quintessence (or, SCDM). For the case of Chaplygin gas model, the trajectories in the $s-r$ plane lie in the region where $s<0$ and $r>1$ [67]. This makes it a powerful tool for effectively distinguishing between various DE models, even when they produce similar expansion histories. However, one can look into the refs. [68, 69, 70, 71, 72, 73, 74, 75, 76, 77], where a detailed statefinder pair analysis for different DE models is comprehensively discussed.
In this work, the present model is scrutinized employing statefinder diagnostic tools. The process involves calculating statefinder parameters with respect to redshift, compare them with that of $\Lambda CDM$, and exploring their behavior at both high and low redshift limits. Using the energy conservation eqs. (15) and (16), we obtain (after some calculations) the energy densities of DM and DE as
@@
$\begin{split}E(z)=\frac{H}{H_{0}}=&(1+z)^{\frac{3}{2}\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}+1\right)}\bigg{[}\Omega_{m0}+\Omega_{\phi 0}~{}(1+z)^{-3\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}-\omega_{\phi}\right)}\\ &\times\left\{1+\frac{\Omega_{m0}\alpha\sqrt{3(1+\omega_{\phi})}}{3\sqrt{3}\Omega_{\phi 0}\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}-\omega_{\phi}\right)}(1+z)^{-3\left(\frac{\alpha}{3}\sqrt{1+\omega_{\phi}}+1\right)}\right\}\bigg{]}^{\frac{1}{2}},\end{split}$ (62)
@@
$\begin{split}r(z)=&\frac{1}{2}\bigg{[}\frac{\Omega_{m0}(z+1)^{-2\alpha\sqrt{\omega_{\phi}+1}+3\omega_{\phi}-3}\left(\Omega_{\phi 0}(z+1)^{\alpha\left(-\sqrt{\omega_{\phi}+1}\right)}\left(9\omega_{\phi}^{2}\Psi\left(3\omega_{\phi}-\alpha\left(3\sqrt{\omega_{\phi}+1}-\alpha\right)\right)+\Theta\right)+\chi\right)}{\Xi^{2}\left(3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}\right)}\\ &+\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+1\right)\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+2\right)\bigg{]},\end{split}$ (64)
@@
$\begin{split}s(z)=&\bigg{[}\frac{\Omega_{m0}(z+1)^{-2\alpha\sqrt{\omega_{\phi}+1}+3\omega_{\phi}-3}\left(\Omega_{\phi 0}(z+1)^{\alpha\left(-\sqrt{\omega_{\phi}+1}\right)}\left(9\omega_{\phi}^{2}\Psi\left(3\omega_{\phi}-\alpha\left(3\sqrt{\omega_{\phi}+1}-\alpha\right)\right)+\Theta\right)+\chi\right)}{\Xi^{2}\left(3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}\right)}\\ &+\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+1\right)\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}+2\right)-2\bigg{]}\div{3\left(\alpha\sqrt{\omega_{\phi}+1}+\frac{\Delta}{\Xi}\right)}\end{split}$ (65)
@@
$\Delta=$ $(z+1)^{-2\alpha\sqrt{\omega_{\phi}+1}+3\omega_{\phi}-3}\bigg{(}-\Omega_{m0}\alpha\sqrt{\omega_{\phi}+1}+\frac{\Omega_{m0}\alpha\left(\alpha+\alpha\omega_{\phi}+3\sqrt{\omega_{\phi}+1}\right)}{3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}}+\Omega_{\phi 0}\left(3\omega_{\phi}-\alpha\sqrt{\omega_{\phi}+1}\right)$
×$\times(z+1)^{\alpha\sqrt{\omega_{\phi}+1}+3}\bigg{)},$
@@
$\chi=$ $-\Omega_{m0}\alpha\left(3\omega_{\phi}^{2}\left(3\sqrt{\omega_{\phi}+1}-4\alpha\right)+2\left(2\alpha^{2}-9\right)\sqrt{\omega_{\phi}+1}\omega_{\phi}+4\alpha\left(\alpha\sqrt{\omega_{\phi}+1}+3\right)+9\sqrt{\omega_{\phi}+1}\right),\mbox{and}~{}$
@@
$-\alpha\left(\alpha\left(\alpha\sqrt{\omega_{\phi}+1}((1+z)^{3\omega_{\phi}}+\Psi)+6(1+z)^{3\omega_{\phi}}\right)+9\sqrt{\omega_{\phi}+1}(1+z)^{3\omega_{\phi}}\right).$
The evolutionary trajectories of the statefinder pair in the $s$-$r$ plane are shown in figures 7 for different choices of the model parameter $\alpha$ and $\omega_{\phi}=-0.95$. The evolutionary trajectories of statefinder pair $(s,r)$ of our model start its evolution along the direction of $s$ increasing and $r$ decreasing and with time, trajectories pass through the $\Lambda CDM$ fixed point $(s=0,r=1)$. After making a swirl, these lies in the region $(s<0,r>1)$ in the future. Also for different values of the model parameter $\alpha$, we found different trajectories and present values of the statefinder pair $(s_{0},r_{0})$ as shown by the colored dots. This scenario shows that $\alpha$ affects on the evolutionary trajectories in the $s$-$r$ plane. Moreover, in figure 8, we obtain same evolutionary curve in $s$-$r$ plane for different values of the model parameter $\alpha$ and $\omega_{\phi}=-1$. In contrary to figure 7, we found that $\alpha$ does not affect on the evolutionary trajectories in the $s$-$r$ plane for the case $\omega_{\phi}=-1$. Furthermore, we have also shown the evolutionary trajectories of another statefinder pair $(q,r)$ for the model in figures 7 and 8 . From figure 7, we found that for $\alpha=0.000001$, the evolutionary curve of statefinder pair $(q,r)$ of the model starts from the $SCDM$ $(r=1,q=0.5)$ in the past while the other evolution curves for the model have a smaller deviation from this fixed point and finally all the curves reach above the de Sitter expansion (SS) $(r=1,q=-1)$ in the future. We also found from figure 8 that $\alpha$ has no effect on the evolutionary trajectories in the $q$-$r$ plane for $\omega_{\phi}=-1$. Moreover, figures 7, 8 clearly indicate that the evolution of $q$ shows a smooth signature flip from its positive value regime to negative value regime in the $q$-$r$ plane. Hence, from the statefinder diagnostics analysis, the present model explains the late-time accelerated universe and also the transition from the early decelerated phase ($q>0$) to the current accelerated phase ($q<0$).